OpenStudy (anonymous):

derivative find dy/dx y=(x+1)(x+2)/(x-1)(x-2)

OpenStudy (solomonzelman):

use (f g) ' = f ' g + g ' f

OpenStudy (anonymous):

i cant do it

OpenStudy (solomonzelman):

tell me the derivative of each: 1) x+1 2) x+2 3) x-1 4) x-2

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

you can do like this also. \[\ln y=\ln \frac{ \left( x+1 \right)\left( x+2 \right) }{ \left( x-1 \right)\left( x-2 \right)}\] \[\ln y=\ln \left\{ \left( x+1 \right)\left( x+2 \right) \right\}-\ln \left\{ \left( x-1 \right)\left( x-2 \right) \right\}\] \[\ln y=\ln \left( x+1 \right)+\ln \left( x+2 \right)-\ln \left( x-1 \right)-\ln \left( x-2 \right)\] diff. w.r.t. x , \[\frac{ 1 }{ y }\frac{ dy }{dx }=\frac{ 1 }{x+1 }+\frac{ 1 }{ x+2 }-\frac{ 1 }{x-1 }-\frac{ 1 }{ x-2 }\] \[\frac{ dy }{dx }=y \left\{ \frac{ 1 }{ x+1 }+\frac{ 1 }{ x+2 }-\frac{ 1 }{ x-1 }-\frac{ 1 }{x-2 } \right\}\] \[\frac{ dy }{dx }=\frac{ \left( x+1 \right)\left( x+2 \right) }{ \left( x-1 \right)\left( x-2\right) }\left\{ \frac{ 1 }{x+1 }+\frac{ 1 }{x+2 }-\frac{ 1 }{x-1 }-\frac{ 1 }{x-2 } \right\}\]